OpenStudy (anonymous):

Solve for y: 2/3y + 2 = 6

OpenStudy (anonymous):

2y/ 3 = 6 2y = 18 -> y = 18/2 = 9

OpenStudy (anonymous):

I suppose you mean 2y/3 not 2/(3y)

OpenStudy (anonymous):

2/3 as in the fraction

OpenStudy (anonymous):

is it (2/3)y or (2/3y)?????

OpenStudy (anonymous):

2 over 3 y

OpenStudy (anonymous):